What's the first step to 5x+5=2/10x+29 ?
1
Expert's answer
2011-03-02T05:42:30-0500
The first simplest step is to combine similar terms
5x+5=2/10x+29 |-5
5x = 2/10x + 24 | *10x
50x2 = 2 + 240x
50x2 -240x -2 =0 | 2
25x2 -120x - 1 = 0
D = 602 + 25 = 3625
x1 = (60 + 5√145)/25 = (12 + √145)/5
x2 = (60 - 5√145)/25 = (12 - √145)/5